OpenStudy (anonymous):

Write the equation of the line that passes through (2, 5) with a slope of 3 in point-slope form. y – 2 = 3(x – 5) y + 8 = 3(x + 5) y + 5 = 3(x + 2) y – 5 = 3(x – 2)

OpenStudy (helder_edwin):

the equatio for such a line is \[ \large y-y_0=m(x-x_0) \] you have \(m=3\) and \((x_0,y_0)\=(2,5))

OpenStudy (helder_edwin):

sorry \[ \large (x_0,y_0)=(2,5) \]

OpenStudy (anonymous):

I'm sorry, What??

OpenStudy (helder_edwin):

the equation u r supposed to write has the form \[ \large y-y_0=m(x-x_0) \] u know this, right??

OpenStudy (anonymous):

No sir..

OpenStudy (helder_edwin):

OK. but how come u have to do this???

OpenStudy (helder_edwin):

anyway \[ \large y-5=3(x-2) \]

OpenStudy (anonymous):

It's a pre test :p I'm sorry about that